# Video[Video][video]

### v2

@@ Abstract
Correlative computational microscopy is accelerating the mapping of dynamic biological systems by integrating morphological and molecular measurements across spatial scales, from organelles to entire organisms. Visualization, measurement, and prediction of interactions among the components of biological systems can be accelerated by generalist computational imaging frameworks that relax the trade-offs imposed by multiplex dynamic imaging. This work reports a generalist framework for wave optical imaging of the architectural order (waveOrder) among biomolecules for encoding and decoding multiple specimen properties from a minimal set of acquired channels, with or without fluorescent labels. waveOrder expresses material properties in terms of elegant physically motivated basis vectors directly interpretable as phase, absorption, birefringence, diattenuation, and fluorophore density; and it expresses image data in terms of directly measurable Stokes parameters. We report a corresponding multi-channel reconstruction algorithm to recover specimen properties in multiple contrast modes. With this framework, we implement multiple 3D computational microscopy methods, including quantitative phase imaging, quantitative label-free imaging with phase and polarization, and fluorescence deconvolution imaging, across scales ranging from organelles to whole zebrafish. These advances are available via an extensible open-source computational imaging library, waveOrder, and a napari plugin, recOrder.
journal: opticajournal
@@ 1 Introduction
Biological functions emerge from the interaction of many components that span length scales, i.e., biomolecules, organelles, cells, tissues, and organs. Correlative imaging of these components’ physical and molecular properties is a growing area of microscopy. Computational imaging methods that optically encode multiple physical and molecular properties and decode them computationally are particularly promising as they relax the trade-offs between spatial resolution, temporal resolution, number of channels, field of view, and sample health that constrain multiplex dynamic imaging.
Correlative label-free and fluorescence imaging "fills the vacuum" of observing a few fluorescent molecules in complex biological environments. These approaches enable measurement of conserved physical properties of cellular compartments [1] and enable dynamic high-throughput imaging for mapping responses of multiple organelles to complex perturbations [2]. Such correlative datasets have enabled the development of neural networks that virtually stain molecular labels from label-free contrast modes [3, 4, 5], further improving the ability to analyze interactions among organelles and cells. Light-sheet fluorescence microscopy has long been used to measure the dynamic interactions of cells [6] and organelles [7, 2]. At scales larger than organelles, recent methods for spatial mapping of gene expression [8], protein distribution [9], and gene perturbations [10] use multi-channel imaging systems that encode multiple molecular species in the imaging data. At the finer scale of biomolecular complexes, cryo-CLEM (correlative light and electron microscopy) reveals the structural basis of the biomolecular function [11, 12].
The microscopy methods used by the above technologies can be modeled via linear image formation models, even though they rely on diverse light-matter interactions. The visualization, measurement, and analysis of specimen properties in these biological studies can be improved by developing a computational imaging framework that reconstructs specimen properties from the acquired multi-channel imaging data. This paper describes a computational imaging framework that addresses this need.
2 Background and prior work
@@
Many works describe microscopic image formation for individual contrast modes including fluorescence contrast [13, 14], phase and absorption contrast [15, 16], and polarization-resolved birefringence contrast [17, 18]. Recently, multiple papers have developed computational imaging methods for integrative measurements of specimen properties, but with models specific to label-free contrast [3, 19, 20, 21, 22] or fluorescence contrast [23, 24]. The waveOrder framework unifies and extends these models to multi-contrast and multi-channel imaging systems that include scalar and polarization-resolved imaging with or without fluorescence labels.
@@
The most general image formation models account for statistical fluctuations in the electric field of light, i.e., coherence, and how the coherence is modulated due to propagation or interaction with matter. Modeling coherence requires bilinear functions (functions whose output depends on pairs of points in the imaging path) and propagation of second-order statistics [25, 15, 26, 27, 28, 29]. We start with this approach before restricting our attention to imaging problems that can be modeled with linear functions. Specifically, we consider spatially incoherent fluorescence samples and label-free imaging systems with a spatially incoherent source.
@@
Several existing frameworks provide some, but not all, of the features of waveOrder. Deconvolution libraries typically focus on fluorescence deconvolution [30, 31], limiting their value in multi-channel correlative settings. Differentiable microscopy libraries [32, 33] flexibly model a wide variety of imaging systems to enable new designs and reconstructions, but they don’t often provide generalist inverse algorithms applicable to diverse contrast mechanisms. waveOrder prioritizes linear reconstruction models for the most widely used computational microscopy contrast methods, facilitating broad applications.
@@
We also find that many reconstruction implementations are not used broadly for biological research because they are not reproducible or easy to access. waveOrder is an open-source and actively developed project that attempts to maximize the usability and impact of linear computational imaging methods by unifying and maintaining the most valuable reconstruction techniques in biological microscopy. waveOrder leverages PyTorch for cross-platform, high-performance analysis of large datasets and integration with learned computer vision models.

### v3

@@ Abstract
Correlative computational microscopy can accelerate imaging and modeling of cellular dynamics by relaxing trade-offs inherent to dynamic imaging. Existing computational microscopy frameworks are either specialized or overly generic, limiting use to fixed configurations or domain experts. We introduce WaveOrder, a generalist wave-optical framework for imaging the architectural order of biomolecules. WaveOrder reconstructs diverse specimen properties from multi-channel acquisitions, with or without fluorescence. It provides a unified representation of linear optical properties and differentiable physics-based image formation models spanning widefield, confocal, light-sheet, and oblique label-free geometries. WaveOrder uses physics-informed ML to auto-tune model parameters and solve blind shift-variant restoration problems. This open-source, PyTorch-based framework enables scalable quantitative imaging across scales from organelles to adult zebrafish, and improves restoration of cellular structures in high-throughput experiments. We validate WaveOrder on diverse imaging applications, demonstrating its ability to recover biomolecular structure beyond the limits of existing approaches.
{refsegment}
@@ 1 Introduction
Biological functions emerge from interactions among components that span length scales, including biomolecules, organelles, cells, tissues, and organs. Capturing the composition and organization of these components therefore requires correlative, high-throughput imaging across spatial, temporal, and molecular dimensions. However, conventional imaging systems are constrained by fundamental trade-offs between spatial resolution, temporal resolution, number of channels, field of view, and sample health, which limit their ability to probe biological processes at scale. Computational imaging methods, which can optically encode multiple physical and molecular properties and decode them computationally, offer a promising route to relaxing these trade-offs. To fully realize this potential, there is a need for a flexible computational framework that supports quantitative reconstruction of biomolecular distributions across diverse imaging configurations at high throughput. This paper introduces a computational imaging framework designed to meet this need.
A generalist computational imaging framework does more than relax tradeoffs imposed by dynamic imaging, it enables correlative microscopy by combining complementary contrast mechanisms through a common physical model. Correlative label-free and fluorescence imaging bridges dense structural information with sparse molecular labeling, linking physical properties of cellular compartments to molecular phenotypes [schlusler_correlative_2022] and supporting downstream tasks including virtual staining [guo_revealing_2020, gomez-de-mariscal_harnessing_2024, liu_robust_2025]. Dynamic correlative imaging extends these benefits to time-resolved studies of cellular responses across perturbations [ivanov_mantis_2024]. Similar strategies appear across modalities: structural and functional OCT (optical coherence tomography) combines scattering, elastography, spectroscopy, and multiphoton excitation to monitor cell dynamics and mechanics in engineered tissues [liang_imaging_2009]; multi-modal label-free metabolic imaging combines two-photon fluorescence, lifetime, and spectral readouts to quantify redox state and mitochondrial function in 3D brain models [zhang_multi-modal_2025]; and at finer scales, cryo-CLEM (correlative light and electron microscopy) combines molecular specificity from fluorescence with ultrastructure from EM to reveal mechanisms inaccessible to any single modality [serwas_getting_2021, pierson_recent_2024]. These examples highlight a consistent need for unified, physics-based frameworks capable of modeling and inverting diverse contrast mechanisms so heterogeneous imaging data can be compared and interpreted quantitatively.
@@
Many works describe microscopic image formation for individual contrast modes including fluorescence contrast [backer_extending_2014, chandler_spatio-angular_2019], phase and absorption contrast [streibl_three-dimensional_1985, bao_quantitative_2016], and polarization-resolved birefringence contrast [torok_imaging_2000, oldenbourg_point-spread_2000]. Several recent works have developed computational imaging methods for multi-channel measurements of specimen properties, but with models specific to label-free contrast [guo_revealing_2020, saba_polarization-sensitive_2021, shin_tomographic_2022, yeh_permittivity_2024, jang_super-resolution_2023] or fluorescence contrast [zhang_six-dimensional_2023, chandler_volumetric_2025]. The WaveOrder framework unifies and extends these models to multi-contrast and multi-channel imaging systems that include scalar and polarization-resolved imaging, with or without fluorescence labels [^1] In this paper, multi-contrast implies different light-matter interactions, e.g., label-free and fluorescence contrast, while multi-channel implies channels acquired with a given contrast, e.g., multiple fluorophores or label-free phase and polarization microscopy..
@@
The most general image formation models account for statistical fluctuations in the electric field of light, i.e., coherence, and how coherence is modulated due to propagation or interaction with matter. Modeling coherence requires bilinear functions (functions whose output depends on pairs of points in the imaging path) and propagation of second-order statistics [streibl_three-dimensional_1985, hopkins_concept_1951, barrett_foundations_2004, mehta_partially_2018, sheppard_partially_2018, wolf_introduction_2007]. We start with this approach before restricting our attention to imaging problems that can be modeled with linear functions. Specifically, we consider spatially incoherent fluorescence samples and label-free imaging systems with a spatially incoherent source. The optical transformation of the specimen’s properties in many widely used imaging systems, such as widefield, confocal, light-sheet, quantitative phase imaging, and quantitative polarization imaging, are well approximated by linear image formation models.
@@
Several existing frameworks provide some, but not all, of the features of WaveOrder (Ext. Data Table 5). Deconvolution libraries typically focus on fluorescence deconvolution [bruce_real-time_2013, sage_deconvolutionlab2_2017, li_incorporating_2022, wernersson_deconwolf_2024, alshaabi_fourier-based_2025, kohli_ring_2025], limiting their value in multi-contrast correlative settings. Differentiable microscopy libraries [herath_differentiable_2022, deb_chromatix_2025, tachella_deepinverse_2025] flexibly model a wide variety of imaging systems to enable new designs and reconstructions, but they typically require a domain expert to match them to a specific application. WaveOrder prioritizes linear reconstruction models for the most widely used computational microscopy contrast methods, facilitating broad applications.
Although deconvolution methods are widely used in microscopy, in practice it is often unclear what point-spread or transfer function to use, making reconstruction a blind deconvolution problem. Blind-deconvolution is even more relevant for high-throughput or large-scale imaging experiments where contrast is shift-variant, as illumination, aberrations, and depth-dependent distortions vary across large fields of view. Classical blind deconvolution has been extensively applied to natural images [levin_understanding_2009], and learning-based approaches are increasingly applied to microscopy data [gilton_model_2021, kang_coordinate-based_2024, alshaabi_fourier-based_2025, volpe_roadmap_2025]. WaveOrder builds on these developments by contributing a physics-guided differentiable computational graph that reduces the dimension of the blind search space and enables auto-tuning of imaging parameters directly from data.
@@
We also find that many reconstruction implementations are not used broadly for biological research because they are not reproducible or easy to access. WaveOrder provides an open-source implementation of linear computational imaging methods that unifies widely used models and reconstruction algorithms within a common framework. WaveOrder leverages PyTorch for cross-platform, high-performance analysis of large datasets and integration with learned computer vision models.
@@
  • a unified and elegant mathematical representation of material properties, illumination, scattering, and detection across a broad range of microscopy imaging systems as a computational graph;
  • a demonstration of reconstructions of biological samples across length scales from organelles to organisms, including auto-tuning of reconstruction parameters with shift-variant data-dependent image quality loss;